# Online Estimation with Rolling Validation: Adaptive Nonparametric Estimation with Streaming Data

### v3

@@ 6.3 Nonparametric SGD: Sieve-type
In this section we study the stability of a genuine online nonparametric estimator: the sieve-SGD (6), combining the treatments of the batch sieve estimators and online parametric SGD estimators. Recall that the estimator $\hat{f}_{i}$ is a weighted linear combination of $J_{i}$ pre-specified basis functions
@@
Very similar to the parametric correspondence, the coefficient updating rule for sieve-SGD is:
@@
$i^{2-4\omega\zeta}E\left[\left\{\nabla_{j}\hat{f}_{i}\left(X\right)\right\}^{2}\mid F^{j}\right]\leq C\quad\text{ almost surely},$
@@
The uniform bound on $X$ and/or $Y$ in Theorem 6.1-6.3 cannot be dropped when establishing the uniform bounds on the stability terms. This is more evident from examining (26) where the calculation is very straightforward. However, as we discussed in Remark 4.4, the almost sure bound on stability is not always necessary. When the estimator converges in a stronger mode to $f_{0}$, such as in $\|\cdot\|_{\infty}$, it can be relaxed to $E\left[\left\{\nabla_{j}\hat{f}_{i}(X)\right\}^{2}\right]\lesssim i^{-2b}$. To establish the bounds on unconditional expectations of the stability terms, we can follow the same treatment but under milder moment assumptions such as $E[\|X\|^{2}]\leq R$ and $E[(Y-f_{0}(X))^{2}\mid X]\leq M$.
@@ 7 Numerical Examples
Figure 7.1: Model selection results of Example 7.1. (A) the true MSE of the candidate models; (B)-(D), average rankings of the models at different sample sizes over 500 repetitions, according to wRV, with weight exponent $\xi=0,1,2$, respectively.
In Figure 7.1, we present averaged simulation results from 500 repetitions. Panel A shows the true distance between the estimators and the underlying regression function, which is a piece of information not available in practice. Panels B-D show the average ranking of the four models, based on rolling statistics wRV with different weight exponent $\xi\in\{0,1,2\}$. A smaller ranking value corresponds to a model preferred by wRV. In Section H.2, we also provide supplementary results using larger $\xi\in\{3,4\}$ and extending the sample size to $10^{6}$ to demonstrate selection consistency for $\xi>0$. A larger $\xi$ can decrease the detection delay as discussed in Section 3.1. Comparing Figure 7.1 and Figure H.1, we can see $\xi\in\{2,3,4\}$ gives very similar choices (measured by the location of the trajectory lines’ crosses), slightly different from $\xi=1$. We do not recommend using $\xi=0$ (Figure 7.1, B) mainly due to its insensitivity to the change of model ranking.
@@
Figure 7.2: Model selection results of Example 2. (A) the true MSE of $8$ models; (B)-(D), average ranking of the models at different sample size, according to wRV. The RV weight exponents are $\xi=0,1,2$, respectively. Averaged over 500 repeats.
@@ 8 Discussion
So far, we have focused on the scenario where data points are generated from an unknown but static distribution. However, the analysis framework presented in this paper can be generalized to accommodate time-varying distributions. We are interested in adapting the procedures when $f_{0}$ experiences, abrupt or incremental, concept drifts. Some variants of the wRV method can provide valuable insights and practical tools for tracking the time-varying model selection target.
@@
We first present the proof of comparing two estimators that are different in convergence rates, that is, under the assumption that $a_{g}>a_{h}$. Denote $H_{n}=\sum_{i=1}^{n}i^{\xi}E[r(h_{i-1})]$. For any $\delta\in(0,1/4)$:
pr( correct selection )$\operatorname{pr}(\text{ correct selection })$ (34)
$=\operatorname{pr}\left(\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\geq\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\right)$
$\geq\operatorname{pr}\left(\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\geq(1-\delta)H_{n}\right.\text{ and }$
@@
$=1-\operatorname{pr}\left(\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\leq(1-\delta)H_{n}\right.\text{ or }$
@@
$\geq 1-\operatorname{pr}\left(\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\leq(1-\delta)H_{n}\right)$
$\quad-\operatorname{pr}\left(\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\geq(1-\delta)H_{n}\right).$
@@
$\operatorname{pr}\left(\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\leq(1-\delta)H_{n}\right)$ (35)
$\leq\operatorname{pr}\left(\left|\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}-H_{n}\right|\geq\delta H_{n}\right)$
$\stackrel{{\scriptstyle(I)}}{{\leq}}\operatorname{pr}\left(\left|\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}-H_{n}\right|\geq\delta n^{1-a_{h}+\xi}\tilde{M}_{h}/2\right)$
@@
$\liminf_{i\rightarrow\infty}i^{a_{h}}E[r(h_{i-1})]$ $=M_{h}$ (36)
$\Rightarrow\liminf_{n\rightarrow\infty}n^{a_{h}-1-\xi}\sum_{i=1}^{n}i^{\xi}E[r(h_{i-1})]$ $\geq M_{h}/(1+\xi-a_{h})=:\tilde{M}_{h}$
@@
To bound $\operatorname{pr}\left(\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\geq(1-\delta)H_{n}\right)$, we denote $G_{n}=\sum_{i=1}^{n}i^{\xi}E[r(g_{i-1})]$. Then we have:
$\operatorname{pr}\left(\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\geq(1-\delta)H_{n}\right)$ (37)
=$\stackrel{{\scriptstyle(I)}}{{=}}\operatorname{pr}\left(\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}-G_{n}\geq(1-\delta)H_{n}-G_{n}\right)$
@@
$\stackrel{{\scriptstyle(II)}}{{\leq}}\operatorname{pr}\left(\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}-G_{n}\geq(1-2\delta)H_{n}\right)$
@@
$\stackrel{{\scriptstyle(III)}}{{\leq}}\operatorname{pr}\left(\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}-G_{n}\geq(1-2\delta)n^{1-a_{h}+\xi}\tilde{M}_{h}/2\right)$
@@
$\operatorname{pr}\left(\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\geq(1-\delta)H_{n}\right)$ (38)
$\leq\operatorname{pr}\left(\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\geq(1-\delta)^{2}n^{1-a_{h}+\xi}\tilde{M}_{h}\right)$
@@
$\stackrel{{\scriptstyle(IV)}}{{\leq}}\operatorname{pr}\left(\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}-G_{n}\geq(1-\delta)^{2}n^{1-a_{h}+\xi}\tilde{M}_{h}-G_{n}\right)$
@@
$\leq\operatorname{pr}\left(\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}-G_{n}\geq\delta(1-\delta)^{2}n^{1-a_{h}+\xi}\tilde{M}_{h}\right)$
@@
$\limsup_{n\rightarrow\infty}n^{a_{h}-1-\xi}\sum_{i=1}^{n}i^{\xi}E[r(g_{i-1})]\leq M_{g}/\left(1+\xi-a_{h}\right)=:\tilde{M}_{g}.$ (39)
@@
$G_{n}\leq(\tilde{M}_{g}+\epsilon)n^{1+\xi-a_{h}}\text{ for any }\epsilon>0.$ (40)
@@
$\lim_{n\rightarrow\infty}n^{a-\xi-1}\sum_{i=1}^{n}i^{\xi}h_{i}=M/(1+\xi-a)=:\tilde{M}.$ (42)
@@
$\left|n^{a-\xi-1}\sum_{i=1}^{n}i^{\xi}h_{i}-\tilde{M}\right|\leq\epsilon_{2}$ (44)
@@
$\operatorname{pr}\left(\left|\sum_{i=1}^{n}i^{\xi}\epsilon_{i}\left\{\hat{f}_{i-1}\left(X_{i}\right)-f_{0}\left(X_{i}\right)\right\}\right|\geq c_{n}\right)\leq 2c_{n}^{-2}\text{var}\left(\sum_{i=1}^{n}i^{\xi}\epsilon_{i}\left\{\hat{f}_{i-1}\left(X_{i}\right)-f_{0}\left(X_{i}\right)\right\}\right)$
@@
=$\stackrel{{\scriptstyle(I)}}{{=}}2c_{n}^{-2}\sum_{i=1}^{n}i^{2\xi}E[\epsilon_{i}^{2}u_{i}]\leq 2\sigma^{2}c_{n}^{-2}\sum_{i=1}^{n}i^{2\xi}E[u_{i}]\stackrel{{\scriptstyle(II)}}{{\rightarrow}}0$
In step $(I)$ we used the noise variables $\epsilon_{i}$ are centered and independent from each other. Recall $\sigma^{2}$ is the bound on the variance of $\epsilon_{i}$ (conditioned on $X_{i}$), for step $(II)$ we combined condition 4.4 and the assumption on sequence $c_{n}$. Note that $\limsup_{i\rightarrow\infty}i^{2a}E\left[\left\{\hat{f}_{i-1}(X)-f_{0}(X)\right\}^{4}\right]=C_{1}$ implies $\limsup_{i\rightarrow\infty}i^{a}E\left[\left\{\hat{f}_{i-1}(X)-f_{0}(X)\right\}^{2}\right]=C_{1}^{1/2}$. ∎
@@
Denote $Z_{i}=u_{i}-E[u_{i}\mid F^{i-1}]$, we have $E[Z_{i}]=0=E[Z_{i}\mid F^{j}]$ for any $j\in[i-1]$. Then applying Chebyshev’s inequality:
$\operatorname{pr}\left(\left|\sum_{i=1}^{n}Z_{i}\right|\geq c_{n}\right)$ $\leq 2c_{n}^{-2}\operatorname{var}\left(\sum_{i=1}^{n}Z_{i}\right)$ (48)
$=2c_{n}^{-2}E\left(\sum_{i=1}^{n}Z_{i}\right)^{2}$
$=2c_{n}^{-2}\left\{\sum_{i=1}^{n}EZ_{i}^{2}+\sum_{i\neq j}E[Z_{i}Z_{j}]\right\}.$
@@
$\operatorname{pr}\left(\left|\sum_{i=1}^{n}Z_{i}\right|\geq c_{n}\right)$ $\leq 2c_{n}^{-2}\sum_{i=1}^{n}EZ_{i}^{2}=2c_{n}^{-2}\sum_{i=1}^{n}E\left(u_{i}-E\left[u_{i}\mid F^{i-1}\right]\right)^{2}$ (50)
$=2c_{n}^{-2}\sum_{i=1}^{n}Eu_{i}^{2}+E[E^{2}[u_{i}\mid F^{i-1}]]-2E[u_{i}E[u_{i}\mid F^{i-1}]]$
$\leq 2c_{n}^{-2}\sum_{i=1}^{n}Eu_{i}^{2}=2c_{n}^{-2}\sum_{i=1}^{n}i^{2\xi}E\left[\left\{\hat{f}_{i-1}\left(X_{i}\right)-f_{0}\left(X_{i}\right)\right\}^{4}\right]\stackrel{{\scriptstyle(I)}}{{\rightarrow}}0$
@@
$c_{n}^{-1}\left|\sum_{i=1}^{n}E[u_{i}|F^{i-1}]-E[u_{i}]\right|\stackrel{{\scriptstyle P}}{{\rightarrow}}0$ (51)
@@
$Ec_{n}^{-2}\left|\sum_{i=1}^{n}E[u_{i}|F^{i-1}]-E[u_{i}]\right|^{2}\rightarrow 0\,,$ (52)
@@
$c_{n}^{-1}\left\|\sum_{i=1}^{n}E[u_{i}|F^{i-1}]-\sum_{i=1}^{n}E[u_{i}]\right\|_{2}=c_{n}^{-1}\left(E\left|\sum_{i=1}^{n}E[u_{i}|F^{i-1}]-\sum_{i=1}^{n}E[u_{i}]\right|^{2}\right)^{1/2}\rightarrow 0$ (53)

### v4

@@ 6.3 Nonparametric SGD: Sieve-type
In this section, we investigate the stability of a genuine online nonparametric estimator: the sieve-SGD (4), which combines the approaches of batch sieve estimators and online parametric SGD estimators. Recall that the estimator $\hat{f}_{i}$ is a weighted linear combination of $J_{i}$ pre-specified basis functions
@@
Very similar to the parametric correspondence, the coefficient update rule for sieve-SGD is:
@@
$i^{2-4\omega\zeta}\mathbb{E}\left[\left\{\nabla_{j}\hat{f}_{i}\left(X\right)\right\}^{2}\mid F^{j}\right]\leq C\quad\text{ almost surely},$
@@
The uniform bound on $X$ and/or $Y$ in Theorem 6.1-6.3 cannot be dropped when establishing the uniform bounds on the stability terms. This is more evident from examining (21) where the calculation is very straightforward. However, as we discussed in Remark 4.4, the almost sure bound on stability is not always necessary. When the estimator converges in a stronger mode to $f_{0}$, such as in $\|\cdot\|_{\infty}$, it can be relaxed to $\mathbb{E}\left[\left\{\nabla_{j}\hat{f}_{i}(X)\right\}^{2}\right]\lesssim i^{-2b}$. To establish the bounds on unconditional expectations of the stability terms, we can follow the same treatment but under milder moment assumptions such as $\mathbb{E}[\|X\|^{2}]\leq R$ and $\mathbb{E}[(Y-f_{0}(X))^{2}\mid X]\leq M$.
@@ 7 Numerical Examples
Figure 7.1: Model selection results of Example 7.1. (A) the true MSE of the candidate models; (B)-(D), average rankings of the models at different sample sizes over 500 repetitions, according to wRV, with weighting exponent $\xi=0,1,2$, respectively.
In Figure 7.1, we present averaged simulation results from 500 repetitions. Panel A shows the true distance between the estimators and the underlying regression function, which is a piece of information not available in practice. Panels B-D show the average ranking of the four models, based on rolling statistics wRV with different weighting exponent $\xi\in\{0,1,2\}$. A smaller ranking value corresponds to a model preferred by wRV. In Section H.2, we also provide supplementary results using larger $\xi\in\{3,4\}$ and extending the sample size to $10^{6}$ to demonstrate selection consistency for $\xi>0$. A larger $\xi$ can decrease the detection delay as discussed in Section 3.1. Comparing Figure 7.1 and Figure H.1, we can see $\xi\in\{2,3,4\}$ gives very similar choices (measured by the location of the trajectory lines’ crosses), slightly different from $\xi=1$. We do not recommend using $\xi=0$ (Figure 7.1, B) mainly due to its insensitivity to the change of model ranking.
@@
Figure 7.2: Model selection results of Example 2. (A) the true MSE of $8$ models; (B)-(D), average ranking of the models at different sample size, according to wRV. The RV weighting exponents are $\xi=0,1,2$, respectively. Averaged over 500 repeats.
@@ 8 Discussion
So far, we have focused on the scenario where data points are generated from an unknown but static distribution. However, the analysis framework presented in this paper can be generalized to accommodate time-varying distributions. We are interested in adapting the procedures when $f_{0}$ experiences abrupt or incremental concept drifts. Some variants of the wRV method can provide valuable insights and practical tools for tracking the time-varying model selection target.
In applications with a large space of hyperparameters or candidate models, the dynamic average method proposed by [37] may be more efficient in searching the candidate space. It would be an interesting future direction to combine our framework with the dynamic averaging scheme and develop theoretically justifiable adaptive model searching methods.
Acknowledgments
We thank the Reviewers and Associate Editors for their constructive comments, which helped to improve this work.
Part of this work was completed while Tianyu Zhang was at Carnegie Mellon University. Tianyu Zhang’s work at Carnegie Mellon University was partially supported by National Institute of Mental Health grant R01MH123184. Jing Lei is partially supported by National Science Foundation grants DMS-2310764 and DMS-2515687.
@@ Acknowledgments
We first present the proof of comparing two estimators that are different in convergence rates, that is, under the assumption that $a_{g}>a_{h}$. Denote $H_{n}=\sum_{i=1}^{n}i^{\xi}\mathbb{E}[r(h_{i-1})]$. For any $\delta\in(0,1/4)$:
( correct selection ) $\mathbb{P}(\text{ correct selection })$ (29)
$=\mathbb{P}\left(\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\geq\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\right)$
$\geq\mathbb{P}\left(\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\geq(1-\delta)H_{n}\right.\text{ and }$
@@
$=1-\mathbb{P}\left(\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\leq(1-\delta)H_{n}\right.\text{ or }$
@@
$\geq 1-\mathbb{P}\left(\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\leq(1-\delta)H_{n}\right)$
$\quad-\mathbb{P}\left(\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\geq(1-\delta)H_{n}\right).$
@@
$\mathbb{P}\left(\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\leq(1-\delta)H_{n}\right)$ (30)
$\leq\mathbb{P}\left(\left|\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}-H_{n}\right|\geq\delta H_{n}\right)$
$\stackrel{{\scriptstyle(I)}}{{\leq}}\mathbb{P}\left(\left|\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}-H_{n}\right|\geq\delta n^{1-a_{h}+\xi}\tilde{M}_{h}/2\right)$
@@
$\liminf_{i\rightarrow\infty}i^{a_{h}}\mathbb{E}[r(h_{i-1})]$ $=M_{h}$ (31)
$\Rightarrow\liminf_{n\rightarrow\infty}n^{a_{h}-1-\xi}\sum_{i=1}^{n}i^{\xi}\mathbb{E}[r(h_{i-1})]$ $\geq M_{h}/(1+\xi-a_{h})=:\tilde{M}_{h}$
@@
To bound $\mathbb{P}\left(\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\geq(1-\delta)H_{n}\right)$, we denote $G_{n}=\sum_{i=1}^{n}i^{\xi}\mathbb{E}[r(g_{i-1})]$. Then we have:
$\mathbb{P}\left(\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\geq(1-\delta)H_{n}\right)$ (32)
=$\stackrel{{\scriptstyle(I)}}{{=}}\mathbb{P}\left(\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}-G_{n}\geq(1-\delta)H_{n}-G_{n}\right)$
@@
$\stackrel{{\scriptstyle(II)}}{{\leq}}\mathbb{P}\left(\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}-G_{n}\geq(1-2\delta)H_{n}\right)$
@@
$\stackrel{{\scriptstyle(III)}}{{\leq}}\mathbb{P}\left(\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}-G_{n}\geq(1-2\delta)n^{1-a_{h}+\xi}\tilde{M}_{h}/2\right)$
@@
$\mathbb{P}\left(\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\geq(1-\delta)H_{n}\right)$ (33)
$\leq\mathbb{P}\left(\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\geq(1-\delta)^{2}n^{1-a_{h}+\xi}\tilde{M}_{h}\right)$
@@
$\stackrel{{\scriptstyle(IV)}}{{\leq}}\mathbb{P}\left(\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}-G_{n}\geq(1-\delta)^{2}n^{1-a_{h}+\xi}\tilde{M}_{h}-G_{n}\right)$
@@
$\leq\mathbb{P}\left(\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}-G_{n}\geq\delta(1-\delta)^{2}n^{1-a_{h}+\xi}\tilde{M}_{h}\right)$
@@
$\limsup_{n\rightarrow\infty}n^{a_{h}-1-\xi}\sum_{i=1}^{n}i^{\xi}\mathbb{E}[r(g_{i-1})]\leq M_{g}/\left(1+\xi-a_{h}\right)=:\tilde{M}_{g}.$
@@
$G_{n}\leq(\tilde{M}_{g}+\epsilon)n^{1+\xi-a_{h}}\text{ for any }\epsilon>0.$
@@
$\lim_{n\rightarrow\infty}n^{a-\xi-1}\sum_{i=1}^{n}i^{\xi}h_{i}=M/(1+\xi-a)=:\tilde{M}.$
@@
$\left|n^{a-\xi-1}\sum_{i=1}^{n}i^{\xi}h_{i}-\tilde{M}\right|\leq\epsilon_{2}$
@@
$\mathbb{P}\left(\left|\sum_{i=1}^{n}i^{\xi}\epsilon_{i}\left\{\hat{f}_{i-1}\left(X_{i}\right)-f_{0}\left(X_{i}\right)\right\}\right|\geq c_{n}\right)\leq 2c_{n}^{-2}\text{var}\left(\sum_{i=1}^{n}i^{\xi}\epsilon_{i}\left\{\hat{f}_{i-1}\left(X_{i}\right)-f_{0}\left(X_{i}\right)\right\}\right)$
@@
=$\stackrel{{\scriptstyle(I)}}{{=}}2c_{n}^{-2}\sum_{i=1}^{n}i^{2\xi}\mathbb{E}[\epsilon_{i}^{2}u_{i}]\leq 2\sigma^{2}c_{n}^{-2}\sum_{i=1}^{n}i^{2\xi}\mathbb{E}[u_{i}]\stackrel{{\scriptstyle(II)}}{{\rightarrow}}0$
In step $(I)$, we used the noise variables $\epsilon_{i}$ that are centered and independent from each other. Recall $\sigma^{2}$ is the bound on the variance of $\epsilon_{i}$ (conditioned on $X_{i}$), for step $(II)$ we combined 4.4 and the assumption on sequence $c_{n}$. Note that $\limsup_{i\rightarrow\infty}i^{2a}\mathbb{E}\left[\left\{\hat{f}_{i-1}(X)-f_{0}(X)\right\}^{4}\right]=C_{1}$ implies $\limsup_{i\rightarrow\infty}i^{a}\mathbb{E}\left[\left\{\hat{f}_{i-1}(X)-f_{0}(X)\right\}^{2}\right]=C_{1}^{1/2}$. ∎
@@
Denote $Z_{i}=u_{i}-\mathbb{E}[u_{i}\mid F^{i-1}]$, we have $\mathbb{E}[Z_{i}]=0=\mathbb{E}[Z_{i}\mid F^{j}]$ for any $j\in[i-1]$. Then, applying Chebyshev’s inequality:
$\mathbb{P}\left(\left|\sum_{i=1}^{n}Z_{i}\right|\geq c_{n}\right)$ $\leq 2c_{n}^{-2}\operatorname{var}\left(\sum_{i=1}^{n}Z_{i}\right)$
$=2c_{n}^{-2}\mathbb{E}\left(\sum_{i=1}^{n}Z_{i}\right)^{2}$
$=2c_{n}^{-2}\left\{\sum_{i=1}^{n}\mathbb{E}Z_{i}^{2}+\sum_{i\neq j}\mathbb{E}[Z_{i}Z_{j}]\right\}.$
@@
$\mathbb{P}\left(\left|\sum_{i=1}^{n}Z_{i}\right|\geq c_{n}\right)$ $\leq 2c_{n}^{-2}\sum_{i=1}^{n}\mathbb{E}Z_{i}^{2}=2c_{n}^{-2}\sum_{i=1}^{n}\mathbb{E}\left(u_{i}-\mathbb{E}\left[u_{i}\mid F^{i-1}\right]\right)^{2}$
$=2c_{n}^{-2}\sum_{i=1}^{n}\mathbb{E}u_{i}^{2}+\mathbb{E}[E^{2}[u_{i}\mid F^{i-1}]]-2\mathbb{E}[u_{i}\mathbb{E}[u_{i}\mid F^{i-1}]]$
$\leq 2c_{n}^{-2}\sum_{i=1}^{n}\mathbb{E}u_{i}^{2}=2c_{n}^{-2}\sum_{i=1}^{n}i^{2\xi}\mathbb{E}\left[\left\{\hat{f}_{i-1}\left(X_{i}\right)-f_{0}\left(X_{i}\right)\right\}^{4}\right]\stackrel{{\scriptstyle(I)}}{{\rightarrow}}0$
@@
$c_{n}^{-1}\left|\sum_{i=1}^{n}\mathbb{E}[u_{i}|F^{i-1}]-\mathbb{E}[u_{i}]\right|\stackrel{{\scriptstyle P}}{{\rightarrow}}0$
@@
$\mathbb{E}c_{n}^{-2}\left|\sum_{i=1}^{n}\mathbb{E}[u_{i}|F^{i-1}]-\mathbb{E}[u_{i}]\right|^{2}\rightarrow 0\,,$
@@
$c_{n}^{-1}\left\|\sum_{i=1}^{n}\mathbb{E}[u_{i}|F^{i-1}]-\sum_{i=1}^{n}\mathbb{E}[u_{i}]\right\|_{2}=c_{n}^{-1}\left(\mathbb{E}\left|\sum_{i=1}^{n}\mathbb{E}[u_{i}|F^{i-1}]-\sum_{i=1}^{n}\mathbb{E}[u_{i}]\right|^{2}\right)^{1/2}\rightarrow 0$ (35)